write an equation of the line that passes through the point (3,-2) with slope of 3
Answer by mananth(16946)

m= 3

Plug value of the slope and point ( 3,-2 ) in
Y = m x + b
-2.00 = 9 + b
b= -2 - 9
b= -11
So the equation will be
Y = 3 x -11
m.ananth@hotmail.ca